OpenStudy (anonymous):

how would you factor out this problem: x^6-2x^5-2x^3+x^4

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

Take x^3 as a common factor first.

OpenStudy (watchmath):

\(x^3(x^3-2x^2-2+x)\) \(x^3(x(x-2)-1(x-2) )\) \(x^3(x-1)(x-2)\)

OpenStudy (watchmath):

ups wrong :)

OpenStudy (watchmath):

\(x^3(x(x-2)+1(x-2))\) \(x^3(x-2)(x+1)\)

OpenStudy (anonymous):

\(x^3(x^3-2x^2+x-2)=x^3(x^2(x-2)+1(x-2)\)\(=x^3(x^2+1)(x-2)\)

OpenStudy (watchmath):

you are right anwar :)